# Modelling reliability of reversible circuits with 2D second-order cellular automata

### v2

@@ 3.2 Damage from improper signals interaction
Despite of possible convenience for consideration of damage propagation due to certain simplification, the point-like defect discussed in Sec. 3.1 could not appear due to normal evolution of RCA. Let us consider some kind of errors due to incorrect interaction between signals such as wrong relative delays or too short intervals between consequent groups of signals.
@@
The similar errors can appear if interval between two consequent groups of signals is too small. For example, such delay should be at least 12 for ‘wide’ swap and 16 for ‘narrow’ swap.
Errors also may depend on particular value of delay and used RCA. Let us consider interaction of two signals approaching in orthogonal directions to the point of intersection of trajectories. The difference of distances between signals and this point can be denoted as $\Delta$. For minimal length of signals ($k=3$) some interaction occurs for $-4\leq\Delta\leq 4$. It explains value of critical range for ‘wide’ swap, because the first signal reach intersection point with additional delay 7 steps in comparison with the second one.
@@
Intersection of signals in swap gates or other kinds of routing can be considered as particular problem of 2D design and resolved by analogue 3D RCA models. However for any non-trivial gate discussed above interaction of signals is necessary process. An elementary interaction relevant already for simple switch gate Figure 2 can be considered to investigate errors due to wrong delays. Such interactions are depicted on Figures 1 – 5 in B.
@@ 3.3 Consequences of damage distribution
For example, already discussed minimum intervals between consequent signals for ‘wide’ and ‘narrow’ swap gates on Figure 1 should be 12 and 16 time steps respectively. For Fredkin gate Figure 3 minimum interval depends on the composition of the signal (representing combinations of zeros and units) in two consequent groups. In the worst case such interval should not be less than 22 time steps. Controlled-NOT gate on Figure 5 can be used for RCA $C_{0}E_{1,2}^{\,I\!I}$, $C_{0}E_{+}^{\,I\!I}$, $C_{0}E_{1-3}^{\,I\!I}$ and $C_{0}E_{1,2o}^{\,I\!I}$ with a relatively short minimum possible interval of 10 time steps between consequent groups of signals.
@@ 4 Discussion and conclusions
Despite of impossibility to use $E_{1,3}^{I\!I}$ for construction of reversible circuits this CA is included in the comparison due to known remarkable properties such as linearity and self-replication of any pattern in $E_{1,3}$ Gard . All CA used here for construction of second-order RCA could be informally treated as some non-linear modifications of the Fredkin replicator $E_{1,3}$.
The analytical formulas Eq. (11) obtained due to linearity of $E_{1,3}^{I\!I}$ also can be extended to $C_{0}E_{1}^{\,I\!I}$ AV14 (and $C_{0}E_{1,3}^{\,I\!I}$) to analyse simple example of damage distribution. Oscillating behaviour of formal dimension ${\mathfrak{D}}$ understandable from such expressions provides some hints about possible issues with instability in damage distribution for considered family of RCA partially confirmed by numerical experiments.
@@
On the other hand, the serious damage from even a single error, along with the difficulty of isolation other elements from its influence, lead to problems with longer reliable work of reversible circuits with many gates partially consistent with conclusion drawn from altentative model with probabilistic RCA RCAnoise .
@@
Only few RCA from rather big families such as CEOT are discussed. However for construction of reversible logic circuits even one of them such as $C_{0}E_{1,2}^{\,I\!I}$ would be enough to demonstrate examples with both conserved and variable number of signals. Other RCA discussed here are useful for analysis of reliability and damage distribution.
For all RCA discussed here computational experiments demonstrate too serious problems for whole circuits even from single errors. Possibility to reduce such effects could be considered as a reasonable research direction.
@@
There is simple measure of complexity for such local rules from CET and CEOT family based on counting number $N_{p}$ of mentioned in Sec. 2.2 pairs $p=(n_{c},n_{e})$ with $f(p)=1$ producing change of central cell. For example, $N_{p}=1$ for $C_{0}E_{1}$ or $C_{0}E_{1}^{\,I\!I}$ and $N_{p}=2$ for $C_{0}E_{1,2}^{\,I\!I}$. For isotropic rule similar number $N_{I}$ could be used, but instead number of different configurations up to rotations and reflections should be used.
For example pair $(c_{c}=0,n_{e}=2)$ includes two different isotropic configurations with neighbouring cells either opposite or with touching corners. The first case is used in definition of $C_{0}E_{1,2o}^{\,I\!I}$. Thus, $N_{I}=1$ for $C_{0}E_{1}^{\,I\!I}$, $N_{I}=2$ for $C_{0}E_{1,2o}^{\,I\!I}$ and $N_{I}=3$ for $C_{0}E_{1,2}^{\,I\!I}$.
@@
Similarly with BBM (billiard ball model of computations) RCA discussed in this work do not use any ‘wires’ for motion between gates. Analogue models with ‘wires’ could be designed using generalization of second-order RCA AV13 implemented in software mentioned earlier CART . Such RCA have five states with additional one for construction of ‘wires’. That could reduce sensitivity to external noise, because all cells around wires are passive and may not participate in damage distribution. Reliability analysis for such RCA with damage distribution only along wires should be discussed elsewhere.
The existence of alternative models with ‘wires’ also justifies earlier informal name ‘wireless ants’ for some RCA discussed here and abbreviations for local rules in software CART such as WlAnt, WlAnt-2, WlAnt-3 and WlAnt-4 for $C_{0}E_{1,2}^{\,I\!I}$, $C_{0}E_{+}^{\,I\!I}$, $C_{0}E_{1-3}^{\,I\!I}$ and $C_{0}E_{1,2o}^{\,I\!I}$ respectively. The local rules for $C_{0}E_{1}^{\,I\!I}$ (WlSnake) and $C_{0}E_{1,3}^{\,I\!I}$ (WlSnake-3) are implemented in the same package. All rules are also converted to so-called ‘tree’ format used for CA simulations in more common software GoL .
@@ Appendix B Modelling of signals interaction
Both $C_{0}E_{1-3}^{\,I\!I}$ and $C_{0}E_{+}^{\,I\!I}$ produce the same damage fronts with ${\mathfrak{D}}_{\infty}=1$, see Figure 4. Despite of some resemblance for $\Delta=0$ of initial time steps shown on Figure 3, ${\mathfrak{D}}_{\infty}=2$ for $C_{0}E_{1,2}^{\,I\!I}$ not only for $|\Delta|=4$, but also for $\Delta=0$ with shape of damage described as diamond-like area with rectangular empty space due to initial impact of $2\times 2$ static blue block.

### v3

@@ 3.2 Damage from improper signals interaction
Despite of possible convenience for consideration of damage propagation due to certain simplification, the point-like defect discussed in Sec. 3.1 could not appear during normal evolution of RCA. Let us consider some kind of errors due to incorrect interaction between signals such as wrong relative delays or too short intervals between consequent groups of signals.
@@
The similar errors can appear if an interval between two consequent groups of signals is too small. For example, the interval should be at least 12 for ‘wide’ swap and 16 for ‘narrow’ swap.
Errors also may depend on the particular value of delay and used RCA. Let us consider the interaction of two signals approaching in orthogonal directions to the point of intersection of trajectories. The difference of distances between signals and this point can be denoted as $\Delta$. For minimal length of signals ($k=3$) some interaction occurs for $-4\leq\Delta\leq 4$. It explains value of critical range for ‘wide’ swap, because the first signal reach intersection point with additional delay 7 steps in comparison with the second one.
@@
Intersection of signals in swap gates or other kinds of routing can be considered as particular problem of 2D design and resolved in 3D RCA models with similar properties. However for any non-trivial gate discussed above interaction of signals is necessary process. An elementary interaction relevant already for simple switch gate Figure 2 can be considered to investigate errors due to wrong delays. Such interactions are depicted on Figures 1 – 5 in B.
@@ 3.3 Consequences of damage distribution
For example, already discussed minimum intervals between consequent signals for ‘wide’ and ‘narrow’ swap gates on Figure 1 should be 12 and 16 time steps respectively. For Fredkin gate on Figure 3 minimum interval depends on the composition of the signal (representing combinations of zeros and units) in two consequent groups. In the worst case such interval should not be less than 22 time steps. Controlled-NOT gate on Figure 5 can be used for RCA $C_{0}E_{1,2}^{\,I\!I}$, $C_{0}E_{+}^{\,I\!I}$, $C_{0}E_{1-3}^{\,I\!I}$ and $C_{0}E_{1,2o}^{\,I\!I}$ with a relatively short minimum possible interval of 10 time steps between consequent groups of signals.
@@ 4 Discussion and conclusions
Despite of impossibility to use $E_{1,3}^{I\!I}$ for construction of reversible circuits this CA is included in the comparison due to known remarkable properties such as linearity and self-replication of any pattern in $E_{1,3}$ [21]. All CA used here for construction of second-order RCA could be informally treated as some non-linear modifications of the Fredkin replicator $E_{1,3}$.
The analytical formulas Eq. (11) obtained due to linearity of $E_{1,3}^{I\!I}$ also can be extended to $C_{0}E_{1}^{\,I\!I}$ [31] (and $C_{0}E_{1,3}^{\,I\!I}$) to analyse simple example of damage distribution. Oscillating behaviour of formal dimension ${\mathfrak{D}}$ understandable from such expressions provides some hints about possible issues with instability in damage distribution for considered family of RCA partially confirmed by numerical experiments.
@@
On the other hand, the serious damage from even a single error, along with the difficulty of isolation other elements from its influence, lead to problems with longer reliable work of reversible circuits with many gates partially consistent with conclusion drawn from altentative model with probabilistic RCA [12].
@@
Only few RCA from rather big families such as CEOT are discussed here. However for construction of reversible logic circuits even one of them such as $C_{0}E_{1,2}^{\,I\!I}$ would be enough to demonstrate examples with both conserved and variable number of signals. Other RCA discussed here are useful for analysis of reliability and damage distribution.
For all RCA presented in this work computational experiments demonstrate too serious problems for whole circuits even from single errors. Possibility to reduce such effects could be considered as a reasonable research direction.
@@
There is simple measure of complexity for such local rules from CET and CEOT family based on counting number $N_{p}$ of mentioned in Sec. 2.2 pairs $p=(n_{c},n_{e})$ with $f(p)=1$ producing change of central cell. For example, $N_{p}=1$ for $C_{0}E_{1}$ or $C_{0}E_{1}^{\,I\!I}$ and $N_{p}=2$ for $C_{0}E_{1,2}^{\,I\!I}$. For isotropic rules similar number $N_{I}$ of different configurations up to rotations and reflections should be used instead.
For example, pair $(n_{c}=0,n_{e}=2)$ includes two different isotropic configurations with two neighboring cells are either opposite or adjacent. The first case (with opposite cells) is used in definition of $C_{0}E_{1,2o}^{\,I\!I}$. Thus, $N_{I}=1$ for $C_{0}E_{1}^{\,I\!I}$, $N_{I}=2$ for $C_{0}E_{1,2o}^{\,I\!I}$ and $N_{I}=3$ for $C_{0}E_{1,2}^{\,I\!I}$.
@@
Similarly with BBM (billiard ball model of computations) RCA discussed in this work do not use any ‘wires’ for motion between gates. Analogue models with ‘wires’ could be designed using generalization of second-order RCA [47] implemented in software mentioned earlier [32]. Such RCA have five states with additional one for construction of ‘wires’. That could reduce sensitivity to external noise, because all cells around wires are passive and may not participate in damage distribution. Reliability analysis for such RCA with damage distribution only along wires should be discussed elsewhere.
The existence of alternative models with ‘wires’ also justifies earlier informal name ‘wireless ants’ for some RCA discussed here and abbreviations for local rules in software [32] such as WlAnt, WlAnt-2, WlAnt-3 and WlAnt-4 for $C_{0}E_{1,2}^{\,I\!I}$, $C_{0}E_{+}^{\,I\!I}$, $C_{0}E_{1-3}^{\,I\!I}$ and $C_{0}E_{1,2o}^{\,I\!I}$ respectively. The local rules for $C_{0}E_{1}^{\,I\!I}$ (WlSnake) and $C_{0}E_{1,3}^{\,I\!I}$ (WlSnake-3) are implemented in the same package. All rules are also converted to so-called ‘tree’ format used for CA simulations in more common software [33].
@@ Appendix B Modelling of signals interaction
Both $C_{0}E_{1-3}^{\,I\!I}$ and $C_{0}E_{+}^{\,I\!I}$ produce the same damage fronts with ${\mathfrak{D}}_{\infty}=1$, see Figure 4. Despite of some resemblance for $\Delta=0$ of initial time steps shown on Figure 3, ${\mathfrak{D}}_{\infty}=2$ for $C_{0}E_{1,2}^{\,I\!I}$ not only for $|\Delta|=4$, but also for $\Delta=0$ with the shape of damage described as a diamond-like area with rectangular empty space due to initial impact of $2\times 2$ static blue block.
Acknowledgements
I would like to thank the anonymous reviewers for their constructive comments and suggestions.